OpenStudy (anonymous):

Jake's band played songs at three different parties. At the first party they played one-fourth as long as they did at the second party, which was one-half as long as the third party. If they played for 7 minutes at the first party, how many minutes did they play at the third party?

OpenStudy (anonymous):

can anyone answer??

OpenStudy (anonymous):

three and one-half minutes fourteen minutes fifty-six minutes one hundred twelve minutes

OpenStudy (anonymous):

there are your answers

OpenStudy (anonymous):

which one is it??

OpenStudy (anonymous):

it's b 14min

OpenStudy (anonymous):

how??

OpenStudy (texaschic101):

I disagree...I think it is 56 minutes

OpenStudy (anonymous):

HOW??

OpenStudy (anonymous):

wait i'm doing

OpenStudy (texaschic101):

a = 1st party b = 2nd party c = 3rd party a = 7 minutes a = 1/4b 7 = 1/4b 7/(1/4) = b 7 * 1/4 = b 28 = b b = 1/2c 28 = 1/2c 28/(1/2) = c 28 * 2/1 = c 56 = c

OpenStudy (anonymous):

its a medal for however gets it right

OpenStudy (texaschic101):

oops...typo...7 * 4/1

OpenStudy (anonymous):

texaschic101 got it right

OpenStudy (shrutipande9):

totally agreed wid @texaschic101

OpenStudy (texaschic101):

sweet :)

OpenStudy (shrutipande9):

:)